No Name University has a successful football team, the Lethargic Leeches, and sells tickets to students, alumni, and the public. Experience has shown that attendance has followed the demand relationship: Q -28,000- 2,000P 4. a. If No Name U charges $7 per ticket, predict attendance (Q). l point c. Now, assume that the current capacity of the football stadium is 10,000 seats, what ticket price should be set for the stadium to be sold-out? l poin
Answer #1

Q = 28,000 - 2,000P

(a) When P = $7,

Q = 28,000 - (2,000 x 7) = 28,000 - 14,000 = 14,000

(b) When Q = 10,000,

10,000 = 28,000 - 2,000P

2,000P = 18,000

P = $9